OpenStudy (anonymous):

find the derivative of y with respect to x, t or theta. (a) e^7 - 10x Answer: -10e^7-10x (b) 8xe^x - 8e^x answer: 8xe^x (c) y= (x^2 -2x+4)e^x answer: (x^2+2)e^x (d) y= sin e^theta^4 answer: (-4theta^3 e^-theta^4) cos e^-theta^4 please show the steps. thank you

jimthompson5910 (jim_thompson5910):

The first one is \[\Large e^{7-10x}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

The derivative of that isn't -10e^7-10x, so there has to be a typo somewhere

OpenStudy (anonymous):

im not sure i think we have to use the u substitution method to find the answer.

jimthompson5910 (jim_thompson5910):

oh wait, is the answer is \[\Large -10e^{7-10x}\] and not \[\Large -10e^{7}-10x\]

OpenStudy (anonymous):

yea its the first one

jimthompson5910 (jim_thompson5910):

this is why parenthesis are useful if 7-10x is in the exponent, then say e^(7-10x)

jimthompson5910 (jim_thompson5910):

ok that makes more sense now

OpenStudy (anonymous):

sorry about that

jimthompson5910 (jim_thompson5910):

let u = 7 - 10x, which means du/dx = -10 this means \[\Large e^{7-10x}\] turns into \[\Large e^{u}\] then derive to get \[\Large e^{u}*\frac{du}{dx}\] \[\Large e^{u}(-10)\] \[\Large -10e^{u}\] \[\Large -10e^{7-10x}\]

jimthompson5910 (jim_thompson5910):

I'm using the chain rule, which is If h(x) = f(g(x)), then h ' (x) = f ' (g(x)) * g ' (x)

OpenStudy (anonymous):

ok i just need help with the other 3 as well

jimthompson5910 (jim_thompson5910):

whats the derivative of 8xe^x

OpenStudy (anonymous):

im not sure

jimthompson5910 (jim_thompson5910):

use the product rule

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

is it 8xe^x

jimthompson5910 (jim_thompson5910):

no, but that's part of it though

jimthompson5910 (jim_thompson5910):

Product Rule: if h(x) = f(x)*g(x), then h ' (x) = f ' (x) * g(x) + f(x) * g ' (x)